# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 5.3 Evolutionary and Search-Based Optimization for Merging Configurations
Since principled importance weighting alone cannot resolve the broader challenge of configuration selection, evolutionary and search-based approaches reformulate merging as an optimization problem over a structured configuration space. These methods enable principled exploration of merge recipes that would be infeasible to discover through exhaustive enumeration or expert intuition alone. Akiba et al. (2019) provide hyperparameter optimi.[truncated] for model merging can be formally characterized as a product of discrete and continuous decision variables. Let $\mathcal{M}=\{M_{1},\ldots,M_{K}\}$ denote a collection of candidate source models. A merge recipe $\theta$ typically consists of a selected model subset $S\subseteq\mathcal{M}$, per-layer mixing coefficients $\alpha^{(l)}\in\Delta^{|S|-1}$ (where $l$ ranges from 1 to 32 or more, depending on model depth), and topological decisions regarding layer permutation or repetition. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathcal{F}(\text{Merge}(\theta))$, where $\mathcal{F}$ represents a fitness function derived from specific validation metrics. In practice, $\mathcal{F}$ is often a composite score from the Open LLM Leaderboard or task-specific benchmarks such as GSM8K for reasoning and HumanEval for coding. This formulation reveals the immense scale of the challenge: a merge of two 70-billion parameter models involves optimizing hundreds of continuous coefficients alongside discrete architectural choices, creating a non-convex landscape rife with local optima.
@@
Evolutionary algorithms (EAs) have emerged as the dominant framework for navigating this space, most notably demonstrated by the landmark work of Akiba et al. (2024). This approach generalizes model merging into two distinct optimization domains: Parameter Space (PS), which optimizes mixing weights (e.g., via SLERP or Task Arithmetic), and Data Flow Space (DFS), which optimizes the inference path by permuting or repeating layers from different models. By applying evolutionary strategies to these spaces, the authors successfully merged a Japanese LLM with a Math-specialized LLM to create a model that achieved state-of-the-art performance on Japanese math reasoning, a capability present in neither parent model individually. Their method employs population-based exploration where ”individuals” (merge recipes) undergo crossover and mutation, allowing the algorithm to discover non-intuitive architectures, such as the ”Franken-merge” structures where layers from different models are interleaved in novel sequences. The FuseLLM framework (Wan et al., 2024a) takes a complementary approach by fusing knowledge from multiple heterogeneous LLMs through lightweight continual training, while FuseChat (Wan et al., 2024a) extends this to chat models specifically.
@@
While Akiba et al. (2024) utilized Tree-structured Parzen Estimators (TPE) and evolutionary heuristics, other population-based strategies like Covariance Matrix Adaptation Evolution Strategy (CMA-ES) offer theoretical advantages for the continuous subspace of mixing coefficients. CMA-ES adaptively learns the covariance structure of high-fitness solutions, concentrating the search in promising regions of the coefficient simplex. However, the computational cost of fitness evaluation remains a bottleneck; typical evolutionary runs reported in the literature may require hundreds of candidate evaluations, consuming significant GPU hours. To mitigate this, practitioners often employ lower-fidelity proxies during the search phase, such as evaluating on small subsets of the validation data (e.g., minibatches of perplexity scores) before verifying top candidates on full benchmarks.
@@
As an alternative to evolutionary methods, Bayesian Optimization (BO) offers a more sample-efficient paradigm for finding optimal merge parameters. BO constructs a probabilistic surrogate model (typically a Gaussian Process) to approximate the fitness function $\mathcal{F}$. By utilizing an acquisition function to balance exploration (sampling uncertain regions) and exploitation (sampling high-performing regions), BO can converge to optimal mixing coefficients with significantly fewer function evaluations than genetic algorithms. However, standard BO struggles with the high-dimensional, mixed-discrete nature of Data Flow Space optimization, making it more suitable for refining mixing coefficients ($\alpha$) once the architectural topology is fixed.
@@
Multi-objective formulations are essential when merged models must satisfy competing criteria, such as balancing safety alignment with reasoning capability. Pareto-based evolutionary algorithms, such as NSGA-II, maintain a population approximating the Pareto frontier, allowing practitioners to select trade-off solutions post-hoc. Such approaches are critical for ”generalist” merging, where a single scalar fitness function fails to capture the degradation in instruction following that often accompanies gains in domain-specific knowledge (Xiao et al., 2023).
@@
Table 7 summarizes the trade-offs between the primary search paradigms employed in automated model merging.
Table 7: Comparison of Search Paradigms for Model Merging
@@
Search-based model merging has shifted the field from manual heuristic design to automated discovery. While evolutionary methods currently lead in discovering complex architectural combinations, the high computational cost of evaluation remains a barrier. Future research will likely converge on hybrid approaches that leverage the structural exploration of evolution with the sample efficiency of Bayesian methods or surrogate-assisted optimization.
@@
Quantitative comparisons on the FusionBench suite (Tang et al., 2024b) reveal that CMA-ES-based search over layer-wise coefficients typically recovers 85–95% of the best oracle merge performance at roughly $10\times$ lower evaluation cost compared to exhaustive grid search, though the absolute gap varies substantially with model scale. For 7B-parameter LLMs, a single evolutionary run requires a non-trivial number of forward-pass evaluations, making it more expensive than data-free methods but still feasible for practitioners with moderate compute budgets.
@@
Evolutionary and search-based methods share a fundamental dependence on the quality of the fitness function. When the evaluation metric is a poor proxy for downstream utility (e.g., perplexity for instruction-following), the discovered merge configurations may be suboptimal in practice. Additionally, the search space grows combinatorially with model depth: a Transformer with $L$ layers and $K$ candidate models admits $\mathcal{O}(K^{L})$ layer-wise configurations, making exhaustive search intractable beyond $K=3$ or $L>40$ without dimensionality reduction techniques such as block-wise grouping. However, these search-based methods implicitly assume that parameters from different models can be meaningfully combined through direct arithmetic operations, an assumption that breaks down when constituent models have diverged substantially during independent fine-tuning, causing their internal representations to become geometrically incompatible despite encoding similar semantic information.
@@ 5.4 Representation-Level Alignment and Cross-Model Matching Techniques
Representation-level alignment and cross-model matching techniques directly address this geometric incompatibility challenge. When neural networks undergo substantial divergence during independent fine-tuning, their internal representations may encode semantically similar information in configurations that defeat simple arithmetic combination. While permutation alignment resolves the discrete symmetry inherent to hidden unit orderings, it assumes that corresponding neurons exist across models in a one-to-one fashion. Representation-level methods relax this assumption by establishing soft correspondences that preserve functional relationships even when model components have undergone more complex transformations.
@@
The theoretical foundation for representation-level alignment rests on the observation that models fine-tuned from shared pretrained weights often learn similar internal representations despite parametric divergence. This phenomenon motivates the application of optimal transport (OT) frameworks to model merging, where the goal is to find a coupling between the weight distributions of source models that minimizes transportation cost while preserving semantic structure. Given two models with weight matrices $W_{A}\in\mathbb{R}^{m\times n}$ and $W_{B}\in\mathbb{R}^{m\times n}$, the optimal transport alignment seeks a permutation or soft assignment matrix $P$ that minimizes $\|W_{A}-PW_{B}\|_{F}^{2}$ subject to constraints encoding valid matchings. Yang et al. (2024b) provide a taxonomy of such alignment strategies, noting that OT-based methods extend naturally to scenarios where exact permutation alignment is insufficient due to representational drift.
@@
Activation-based correspondence discovery offers an alternative paradigm that bypasses direct weight comparison by instead analyzing how models process identical inputs. The core insight is that layers with similar functional roles will produce correlated activation patterns on shared calibration data, even when their weights have diverged substantially. Formally, given activation matrices $A_{A}^{(l)}$ and $A_{B}^{(k)}$ from layers $l$ and $k$ of models $A$ and $B$ respectively, cross-model correspondence can be quantified through centered kernel alignment (CKA) or related similarity metrics. This enables not only same-layer alignment but also the discovery of correspondences between layers at different depths, which proves essential when fine-tuning has induced differential layer-wise adaptation rates.
@@
Cross-model feature matching extends these concepts by employing auxiliary objectives to guide alignment. Probe tasks, linear classifiers trained on intermediate representations to predict semantic properties, provide a mechanism for identifying which components encode functionally equivalent information. Yadav et al. (2024a) emphasize that model collaboration benefits substantially from such representation-level understanding, as it enables principled composition of specialized capabilities. Unlike purely geometric approaches that optimize alignment based on parameter statistics, probe-guided methods incorporate task-relevant semantic information, potentially yielding alignments that better preserve downstream performance.
@@
Handling architectural variations requires representation projection techniques that map activations from one model’s feature space into another’s. When source models differ in hidden dimensions or layer configurations, direct weight averaging becomes undefined. Projection-based approaches learn linear or nonlinear transformations $\phi:\mathbb{R}^{d_{A}}\rightarrow\mathbb{R}^{d_{B}}$ that minimize reconstruction error on calibration activations while preserving task-relevant structure. These techniques connect intimately to neural network stitching, where layers from different models are composed through learned adapter modules. Yadav et al. (2024a) demonstrate that such modular composition strategies can effectively integrate diverse expert capabilities while maintaining computational efficiency through sparse activation patterns.
@@
The relationship between representation-level alignment and neural network stitching reveals a spectrum of integration strategies. Stitching approaches preserve source model components intact while learning minimal connective tissue, whereas alignment-based merging produces unified parameter sets that blend source contributions. This distinction between preserving model structure versus unifying parameters carries significant implications for computational efficiency, memory requirements, and the preservation of specialized capabilities, considerations that become particularly salient when scaling to multiple expert models.
@@ 5.5 Comparative Analysis of Structural Preservation versus Parameter Unification Tradeoffs
These implications become concrete when examining specific deployment scenarios where the structural preservation versus parameter unification dichotomy manifests as practical tradeoffs. Parameter unification methods, including weight averaging, task vector arithmetic, and Fisher-weighted merging, produce a single model with identical inference costs to any source model, making them attractive for resource-constrained deployment. Conversely, structural preservation approaches such as Mixture-of-Experts (MoE) architectures maintain distinct expert pathways, incurring additional memory overhead proportional to the number of integrated experts. Yang et al. (2024b) establish a comprehensive taxonomy that highlights this tradeoff, noting that unified models achieve parameter efficiency at the potential cost of capability interference, while structured approaches preserve expert specialization through architectural partitioning.
@@
The memory-computation tradeoff can be formalized by considering a base model with $N$ parameters being merged from $K$ expert sources. Parameter unification produces a model of size $N$, whereas dense MoE preservation requires $KN$ parameters plus routing overhead $R$. However, sparse activation patterns mitigate inference costs: if only $k\ll K$ experts activate per token, effective computation scales as $kN+R$ rather than $KN$. demonstrates that weight-ensembling MoE approaches can achieve favorable efficiency-capability tradeoffs by dynamically combining expert contributions through learned routing, effectively interpolating between full structural preservation and parameter unification (Yadav et al., 2024a).
@@
Capability retention patterns diverge substantially across these paradigms. Unified merging methods exhibit interference phenomena when source tasks compete for shared parameters, manifesting as degraded performance on individual tasks despite improved average multi-task scores. Akiba et al. (2024) find that model merging transcends simple aggregation, functioning as a transformative process that can yield emergent capabilities (Wei et al., 2022) exceeding source models. Nevertheless, this transformation introduces unpredictability in capability preservation. Structural preservation approaches circumvent direct interference by maintaining task-specific pathways, achieving near-complete retention of individual expert capabilities at the cost of increased architectural complexity.
@@
Flexibility considerations further differentiate these paradigms. Unified models support straightforward continual adaptation through additional fine-tuning or subsequent merging operations, as they conform to standard architectural specifications. Structural approaches require more sophisticated modification protocols: adding new experts necessitates router retraining, while removing experts may disrupt learned routing distributions. Yadav et al. (2024a) emphasize that model collaboration architectures must balance modularity against integration complexity, with structural preservation offering superior compositionality for scenarios requiring frequent expert addition or replacement.
@@
A decision framework for practitioners emerges from synthesizing these considerations. Parameter unification suits deployment-constrained scenarios where inference efficiency is paramount and moderate capability interference is acceptable, typical in general-purpose assistants requiring broad competence across multiple tasks. Structural preservation addresses applications demanding peak performance on specialized tasks without compromise, such as domain-specific enterprise deployments where computational resources are less constrained than accuracy requirements. Hybrid approaches, exemplified by sparse MoE with weight-ensembled experts, occupy an intermediate position suitable for applications requiring scalable expertise integration with controlled computational growth. The optimal choice ultimately depends on the relative prioritization of capability fidelity, deployment efficiency, and architectural flexibility within specific application contexts. Understanding how unified parameter approaches achieve multi-task competence requires examining the mechanisms through which merged models acquire and combine diverse capabilities from their constituent sources.
@@
The structured and information-guided approaches examined in this section represent a methodological shift from treating model merging as purely parameter-space interpolation toward leveraging architectural knowledge, activation patterns, and systematic search to preserve functional capabilities while combining multiple models. These techniques navigate fundamental tradeoffs between maintaining expert specialization through routing mechanisms and achieving parameter efficiency through unification, with evolutionary and representation-alignment methods offering principled frameworks for discovering optimal configurations. With this technical foundation in place, the next section examines where and why such merging proves valuable in practice, organizing applications by their underlying motivations, whether seeking enhanced capabilities, improved alignment properties, or computational efficiency gains.

### 30 Mar 2026 v2

@@ 5.3 Evolutionary and Search-Based Optimization for Merging Configurations
Since principled importance weighting alone cannot resolve the broader challenge of configuration selection, evolutionary and search-based approaches reformulate merging as an optimization problem over a structured configuration space. These methods open the door to principled exploration of merge recipes that would be infeasible to discover through exhaustive enumeration or expert intuition alone. Akiba et al. (2019) provide hyperparameter optimization tools commonly used in this setting. The search space for model merging can be formally characterized as a product of discrete and continuous decision variables. Let $\mathcal{M}=\{M_{1},\ldots,M_{K}\}$ denote a collection of candidate source models. A merge recipe $\theta$ typically consists of a selected model subset $S\subseteq\mathcal{M}$, per-layer mixing coefficients $\alpha^{(l)}\in\Delta^{|S|-1}$ (where $l$ ranges from 1 to 32 or more, depending on model depth), and topological decisions regarding layer permutation or repetition. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathcal{F}(\text{Merge}(\theta))$, where $\mathcal{F}$ represents a fitness function derived from specific validation metrics. In practice, $\mathcal{F}$ is often a composite score from the Open LLM Leaderboard or task-specific benchmarks such as GSM8K for reasoning and HumanEval for coding. This formulation reveals the immense scale of the challenge. A merge of two 70-billion parameter models involves optimizing hundreds of continuous coefficients alongside discrete architectural choices, creating a non-convex landscape rife with local optima.
@@
Evolutionary algorithms (EAs) are the dominant framework for navigating this space, most notably shown by the landmark work of Akiba et al. (2025). This approach generalizes model merging into two distinct optimization domains, namely Parameter Space (PS), which optimizes mixing weights (e.g., via SLERP or Task Arithmetic), and Data Flow Space (DFS), which optimizes the inference path by permuting or repeating layers from different models. By applying CMA-ES (Covariance Matrix Adaptation Evolution Strategy) to optimize continuous mixing weights in PS and evolutionary search over discrete layer permutations in DFS, the authors successfully merged a Japanese LLM with a Math-specialized LLM to create a model that achieved top performance on Japanese math reasoning, a capability present in neither parent model individually. CMA-ES adaptively learns the covariance structure of high-fitness solutions, concentrating the search in promising regions of the coefficient space and allowing the algorithm to discover non-intuitive architectures, such as the “Franken-merge” structures where layers from different models are interleaved in novel sequences. EvoMerge (Jiang, 2024) extends neuroevolution principles specifically for LLM merging, while CycleQD (Kuroki et al., 2025) shows that quality-diversity algorithms can discover diverse merging configurations for agent skill acquisition. Mergenetic (Minut et al., 2025) provides a simple and accessible evolutionary model merging library that democratizes these techniques for practitioners.
@@
While Akiba et al. (2025) demonstrated the effectiveness of CMA-ES for the continuous subspace of mixing coefficients, the computational cost of fitness evaluation remains a bottleneck; typical evolutionary runs reported in the literature may require hundreds of candidate evaluations, consuming significant GPU hours. To mitigate this, practitioners often use lower-fidelity proxies during the search phase, such as evaluating on small subsets of the validation data (e.g., minibatches of perplexity scores) before verifying top candidates on full benchmarks.
@@
As an alternative to evolutionary methods, Bayesian Optimization (BO) offers a more sample-efficient strategy for finding optimal merge parameters. BO constructs a probabilistic surrogate model (typically a Gaussian Process) to approximate the fitness function $\mathcal{F}$. By using an acquisition function to balance exploration (sampling uncertain regions) and exploitation (sampling high-performing regions), BO can converge to optimal mixing coefficients with far fewer function evaluations than genetic algorithms. Jang et al. (2024a) show that Bayesian optimization produces competitive merging quality with markedly reduced evaluation budgets compared to evolutionary approaches, particularly when the search space is restricted to continuous mixing coefficients. SIP-BMM (Chen et al., 2025b) extends this direction by constructing capability-efficiency Pareto sets through Bayesian model merging with structural importance priors. However, standard BO struggles with the high-dimensional, mixed-discrete nature of Data Flow Space optimization, making it more suitable for refining mixing coefficients ($\alpha$) once the architectural topology is fixed.
@@
Multi-objective formulations are essential when merged models must satisfy competing criteria, such as balancing safety alignment with reasoning capability. Pareto-based evolutionary algorithms, such as NSGA-II, maintain a population approximating the Pareto frontier, allowing practitioners to select trade-off solutions post-hoc. Such approaches are critical for ”generalist” merging, where a single scalar fitness function fails to capture the degradation in instruction following that often accompanies gains in domain-specific knowledge (Xiao et al., 2024).
@@
Table 7 summarizes the trade-offs between the primary search paradigms used in automated model merging.
Table 7: Search paradigms for automated model merging configuration discovery. Evolutionary algorithms excel at exploring discrete architectural choices (e.g., layer permutation) but require many fitness evaluations; Bayesian optimization is more sample-efficient for continuous coefficient tuning; random search serves as a scalable baseline. “Variable Handling” indicates support for mixed discrete-continuous search spaces.
@@
Search-based model merging has shifted the field from manual heuristic design to automated discovery. While evolutionary methods currently lead in discovering complex architectural combinations, the high computational cost of evaluation remains a barrier. Future research will likely converge on hybrid approaches that combine the structural exploration of evolution with the sample efficiency of Bayesian methods or surrogate-assisted optimization.
@@
Quantitative comparisons on the FusionBench suite (Tang et al., 2024c) reveal that CMA-ES-based search over layer-wise coefficients typically recovers a large fraction of the best oracle merge performance at much lower evaluation cost compared to exhaustive grid search, though the absolute gap varies with model scale. A multi-fidelity framework (Su and Geiping, 2025) for automated model merging further reduces search costs by hierarchically evaluating candidates at increasing fidelity levels. MERGE3 (Mencattini et al., 2025) makes evolutionary merging feasible on consumer-grade GPUs, showing that search-based approaches need not be restricted to resource-rich settings. Nature-inspired population-based evolution (Zhang et al., 2025f) explores alternative bio-inspired optimization strategies for LLM merging. For 7B-parameter LLMs, a single evolutionary run requires a non-trivial number of forward-pass evaluations, making it more expensive than data-free methods but still feasible for practitioners with moderate compute budgets.
@@
Evolutionary and search-based methods share a fundamental dependence on the quality of the fitness function. When the evaluation metric is a poor proxy for downstream utility (e.g., perplexity for instruction-following), the discovered merge configurations may be suboptimal in practice. Additionally, the search space grows combinatorially with model depth. A Transformer with $L$ layers and $K$ candidate models admits $\mathcal{O}(K^{L})$ layer-wise configurations, making exhaustive search intractable beyond $K=3$ or $L>40$ without dimensionality reduction techniques such as block-wise grouping. However, these search-based methods implicitly assume that parameters from different models can be meaningfully combined through direct arithmetic operations, an assumption that breaks down when constituent models have diverged substantially during independent fine-tuning, causing their internal representations to become geometrically incompatible despite encoding similar semantic information.
@@ 5.4 Representation-Level Alignment and Cross-Model Matching Techniques
Representation-level alignment and cross-model matching techniques directly confront this geometric incompatibility challenge. When neural networks undergo substantial divergence during independent fine-tuning, their internal representations may encode semantically similar information in configurations that defeat simple arithmetic combination. While permutation alignment resolves the discrete symmetry inherent to hidden unit orderings, it assumes that corresponding neurons exist across models in a one-to-one fashion. Representation-level methods relax this assumption by establishing soft correspondences that preserve functional relationships even when model components have undergone more complex transformations.
@@
Representation-level alignment rests on the theoretical observation that models fine-tuned from shared pretrained weights often learn similar internal representations despite parametric divergence. This phenomenon motivates the application of optimal transport (OT) frameworks to model merging, where the goal is to find a coupling between the weight distributions of source models that minimizes transportation cost while preserving semantic structure. Given two models with weight matrices $W_{A}\in\mathbb{R}^{m\times n}$ and $W_{B}\in\mathbb{R}^{m\times n}$, the optimal transport alignment seeks a permutation or soft assignment matrix $P$ that minimizes $\|W_{A}-PW_{B}\|_{F}^{2}$ subject to constraints encoding valid matchings. Yang et al. (2024b) provide a taxonomy of such alignment strategies, noting that OT-based methods extend naturally to scenarios where exact permutation alignment is insufficient due to representational drift.
@@
Activation-based correspondence discovery offers an alternative approach that bypasses direct weight comparison by instead analyzing how models process identical inputs. The core insight is that layers with similar functional roles will produce correlated activation patterns on shared calibration data, even when their weights have diverged substantially. Interestingly, Hendel et al. (2023) show that in-context learning creates implicit task vectors within the representation space, suggesting deep connections between the task vector framework and internal model representations. Neuron alignment with fixed anchors (Li et al., 2026c) proposes establishing correspondence through training-time alignment, providing a complementary approach to post-hoc matching. Formally, given activation matrices $A_{A}^{(l)}$ and $A_{B}^{(k)}$ from layers $l$ and $k$ of models AA and $B$ respectively, cross-model correspondence can be quantified through centered kernel alignment (CKA) or related similarity metrics. This allows not only same-layer alignment but also the discovery of correspondences between layers at different depths, which proves essential when fine-tuning has induced differential layer-wise adaptation rates. Beyond the Permutation Symmetry of Transformers (Zhang et al., 2025a) extends alignment theory by analyzing the role of rotation symmetries specific to Transformer architectures, revealing additional symmetry structures beyond classical permutation invariance that affect merging.
@@
Cross-model feature matching extends these concepts by using auxiliary objectives to guide alignment. Probe tasks, linear classifiers trained on intermediate representations to predict semantic properties, provide a mechanism for identifying which components encode functionally equivalent information. Yadav et al. (2025a) emphasize that model collaboration benefits from such representation-level understanding in concrete ways, as it enables principled composition of specialized capabilities. Unlike purely geometric approaches that optimize alignment based on parameter statistics, probe-guided methods incorporate task-relevant semantic information, potentially yielding alignments that better preserve downstream performance.
@@
Handling architectural variations requires representation projection techniques that map activations from one model’s feature space into another’s. When source models differ in hidden dimensions or layer configurations, direct weight averaging becomes undefined. Projection-based approaches learn linear or nonlinear transformations $\phi:\mathbb{R}^{d_{A}}\rightarrow\mathbb{R}^{d_{B}}$ that minimize reconstruction error on calibration activations while preserving task-relevant structure. These techniques connect intimately to neural network stitching, where layers from different models are composed through learned adapter modules. Such modular composition strategies can effectively integrate diverse expert capabilities while maintaining computational efficiency through sparse activation patterns (Yadav et al., 2025a).
@@
Contrasting representation-level alignment with neural network stitching reveals a spectrum of integration strategies. Stitching approaches preserve source model components intact while learning minimal connective tissue, whereas alignment-based merging produces unified parameter sets that blend source contributions. This distinction between preserving model structure versus unifying parameters carries practical implications for computational efficiency, memory requirements, and the preservation of specialized capabilities, considerations that become particularly salient when scaling to multiple expert models.
@@ 5.5 Comparative Analysis of Structural Preservation versus Parameter Unification Tradeoffs
These implications become concrete when examining specific deployment scenarios where the structural preservation versus parameter unification dichotomy manifests as practical tradeoffs. Parameter unification methods, including weight averaging, task vector arithmetic, and Fisher-weighted merging, produce a single model with identical inference costs to any source model, making them attractive for resource-constrained deployment. Conversely, structural preservation approaches such as Mixture-of-Experts (MoE) architectures maintain distinct expert pathways, incurring additional memory overhead proportional to the number of integrated experts. Yang et al. (2024b) establish a detailed taxonomy that highlights this tradeoff, namely that unified models achieve parameter efficiency at the potential cost of capability interference, whereas structured approaches preserve expert specialization through architectural partitioning.
@@
To formalize the memory-computation tradeoff, consider a base model with NN parameters being merged from $K$ expert sources. Parameter unification produces a model of size NN, whereas dense MoE preservation requires KNKN parameters plus routing overhead $R$. However, sparse activation patterns mitigate inference costs, since only $k\ll K$ experts activate per token, so effective computation scales as $kN+R$ rather than KNKN. Weight-ensembling MoE approaches can strike favorable efficiency-capability tradeoffs by dynamically combining expert contributions through learned routing, effectively interpolating between full structural preservation and parameter unification (Yadav et al., 2025a).
@@
Capability retention patterns diverge markedly across these approaches. Unified merging methods exhibit interference phenomena when source tasks compete for shared parameters, manifesting as degraded performance on individual tasks despite improved average multi-task scores. Akiba et al. (2025) find that model merging transcends simple aggregation, functioning as a transformative process that can yield emergent capabilities (Wei et al., 2022) exceeding source models. Nevertheless, this transformation introduces unpredictability in capability preservation. Structural preservation approaches circumvent direct interference by maintaining task-specific pathways, attaining near-complete retention of individual expert capabilities at the cost of increased architectural complexity.
@@
Flexibility considerations further differentiate these method families. Unified models support straightforward continual adaptation through additional fine-tuning or subsequent merging operations, as they conform to standard architectural specifications. Structural approaches require more complex modification protocols, as adding new experts necessitates router retraining, while removing experts may disrupt learned routing distributions. Yadav et al. (2025a) emphasize that model collaboration architectures must balance modularity against integration complexity, with structural preservation offering superior compositionality for scenarios requiring frequent expert addition or replacement.
@@
A decision guide for practitioners follows from synthesizing these considerations. Parameter unification suits deployment-constrained scenarios where inference efficiency is paramount and moderate capability interference is acceptable, typical in general-purpose assistants requiring broad competence across multiple tasks. Structural preservation suits applications demanding peak performance on specialized tasks without compromise, such as domain-specific enterprise deployments where computational resources are less constrained than accuracy requirements. Hybrid approaches, exemplified by sparse MoE with weight-ensembled experts, occupy an intermediate position suitable for applications requiring scalable expertise integration with controlled computational growth. The optimal choice ultimately depends on the relative prioritization of capability fidelity, deployment efficiency, and architectural flexibility within specific application contexts. Understanding how unified parameter approaches achieve multi-task competence requires examining the mechanisms through which merged models acquire and combine diverse capabilities from their constituent sources.
Collectively, the structured and information-guided approaches examined in this section represent a methodological shift from treating model merging as purely parameter-space interpolation toward exploiting architectural knowledge, activation patterns, and systematic search to preserve functional capabilities while combining multiple models. Recent work continues to advance along several complementary axes, each navigating the central tension between preserving expert specialization and achieving parameter efficiency.
In the MoE construction and compression space, a key challenge is that expert-based merging produces models with high inference costs proportional to the number of experts. Recent methods address this via post hoc compression. Self-MoE (Kang et al., 2025b) achieves compositional LLMs with self-specialized experts, while Sub-MoE (Li et al., 2026a), PuzzleMoE (Zhao et al., 2025), and MergeMoE (Miao et al., 2025) compress large MoE models via subspace, sparse, and output-based expert merging respectively. Channel Merging (Zhang et al., 2025e) preserves specialization for merged experts. The common thread is a two-stage workflow, first constructing a modular MoE from independently fine-tuned models, then compressing it back toward a parameter-efficient form, retaining the routing-based flexibility without the full cost.
For expert routing and selection, the quality of the routing mechanism determines whether the merged model gracefully degrades or catastrophically fails on out-of-distribution inputs. CAMEx (Nguyen et al., 2025b) introduces curvature-aware expert selection that accounts for the local geometry of the loss landscape; Nash Bargaining (Nguyen et al., 2025a) applies game-theoretic principles to balance expert contributions; Local Mixtures of Experts (Bertolissi et al., 2025) supports test-time training via merging; and Efficient Pareto Set Approximation (Tang et al., 2024a) uses MoE-based fusion to approximate the multi-objective Pareto front.
LoRA-specific composition has emerged as a particularly active subfield, driven by the proliferation of parameter-efficient adapters on open model hubs. The challenge here is distinct from full-model merging, since LoRA modules are low-rank and task-specific, so merging them raises questions about rank allocation and subspace compatibility. CoMoL (Cao et al., 2026a) proposes dynamic core space LoRA merging; MeteoRA (Xu et al., 2025c) embeds multiple tasks in LoRA modules; HydraOpt (Ceritli et al., 2025) navigates adapter merging trade-offs; Tensorized Clustered LoRA (Su et al., 2025) reduces multi-task interference; FlyLoRA (Zou et al., 2025) introduces rank-wise mixture-of-experts; IterIS (Chen et al., 2025a) tackles inference-solving alignment; Adaptive LoRA Merge (Miyano and Arase, 2025) combines pruning with merging; Exploring Sparse Adapters (Arnob et al., 2025) scales parameter-efficient expert merging; and Liu et al. (2026) investigate the practical reality of LoRA recycling. A recurring finding is that methods designed for full-model merging do not transfer directly to the LoRA setting, motivating adapter-specific algorithms.
Finally, automated search methods such as AutoMerge (Lu et al., 2026) address the combinatorial explosion of merging configurations by treating recipe discovery as a meta-optimization problem, automating the selection of which models to merge, at which granularity, and with which algorithm.
@@
These techniques collectively navigate the fundamental tradeoff between maintaining expert specialization through routing mechanisms and achieving parameter efficiency through unification. With this technical foundation in place, the next section examines where and why such merging proves valuable in practice, organizing applications by their underlying motivations, whether seeking enhanced capabilities, improved alignment properties, or computational efficiency gains.